FIND THE AREA BETWEEN THE CURVES Y=X³ AND Y=X

Grade:8

4 Answers

PERUMAL
10 years ago
please write the intervals or correct the question once again.
PERUMAL
10 years ago
area=pi r^2 A(x) =x^3-x =x^6-x^2 ANS:3/4 6
Sher Mohammad IIT Delhi
10 years ago
X^3=X
gives X=0, AND X=1


Sher Mohammad,
B.Tech, IIT Delhi
9247710188
10 years ago
no